Grade 12General Physics

For a given Kinetic Energy, which of the following has smallest De- Broglie wave lengt- electron, proton and a8 – particle? ?

sudhanshu
12 Years agoGrade 12

1 Answer

Apoorva Arora
12 Years ago
The de-broglie wavelength is given by
\lambda =\frac{h}{P}
where
P=\sqrt{2mKE}
so
\lambda =\frac{h}{\sqrt{2mKE}}
so for a given KE, the wavelength is inversely proportional to the square root of mass. Since, alpha particle is the heaviest among all, Its wavelength is the shortest.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty